OpenStudy (anonymous):

Which of the following is a polynomial with roots –2, –2i, and 2i ? x^3 – 4x^2 + 4x – 16 x^3 + 4x^2 + 4x + 16 x^3 + 2x^2 + 4x + 8 x^3 – 2x^2 + 4x – 8

OpenStudy (anonymous):

Its the third answer

OpenStudy (anonymous):

u sure? how'd u get it?

OpenStudy (anonymous):

mutiply the roots; i equals sqrt(-1) so when you multiply them together you get -1

OpenStudy (anonymous):

\[(x+2)(x+2i)(x-2i)\] \[(x+2)(x-2\sqrt{-1})(x+2\sqrt{-1})\] \[(x+2)(x^2+2xsqrt{-1}-2x \sqrt{-1}-4(-1))\] \[(x+2)(x^2+4)\] \[x^3+2x^2+4x+8\]